Swapple Inc. is considering a capital project involving a $40,000 investment in machines and a $40,000 investment in working capital. The equipment is estimated to produce a net cash benefit of $12,000 for each of the next five years. The $40,000 investment in working capital will be recovered at the end of five years. The company has a cost of capital of 14%. Given the present value of an annuity of $1 for five periods at 14% is 3.43308 and the present value of $1 at 14% received after 5 periods is 0.51937, what is the net present value of the investment?
Answer #1

Present value of inflows=$12000*Present value of annuity factor(14%,5)+$40,000*Present value of discounting factor(14%,5)

=12000*3.43308+40,000*0.51937

=61971.76

Present value of outflows=(40,000+40,000)=$80,000

NPV=Present value of inflows-Present value of outflows

=61971.76-80,000

=(18028.24)(Negative).